OpenStudy (anonymous):

what is the quadratic equation for x^2-8x+10=-4

OpenStudy (texaschic101):

The quadratic equation is set up like this : ax + by + c = 0 so lets set up x^2 - 8x + 10 = -4 into the quadratic equation. first, lets add 4 to both sides x^2 - 8x + 10 + 4 = -4 + 4 --- combine like terms x^2 - 8x + 14 = 0 <== there ya go :)